OpenStudy (anonymous):

In a circle, the circumference and diameter vary directly. Which of the following equations would allow you to find the diameter of a circle with a circumference of 154 if you know that in a second circle the diameter is 14 when the circumference is 44? A.154d = (14)(44) B.154/d = 44/14 C.14d = 154/44

OpenStudy (anonymous):

john help me

OpenStudy (johnweldon1993):

lol I just started reading it hang on haha

OpenStudy (anonymous):

ok :) i wait

OpenStudy (johnweldon1993):

ugh let me do that all over again!!! lol too many typos!!

OpenStudy (anonymous):

If x varies directly as y, and x = 7.5 when y = 10, find x when y = 4. A.3 B.5 1/3 C.18.75

OpenStudy (anonymous):

I UNDERSTANED

OpenStudy (johnweldon1993):

haha are you sure? lol I'm gonna do it again anyways :P Varying directly means \[C = kD\] Plug in what we know and solve for k: 44=k(14) k=44/14 so now that we know what 'k' equals...and we want to solve for d... \[154=(\frac{ 44 }{ 14 })(D)\] Divide both sides by D to get... \[\frac{ 154 }{ D }=\frac{ 44 }{ 14 }\] So that looks like B to me... there much better :)

OpenStudy (anonymous):

If x varies directly as y, and x = 7.5 when y = 10, find x when y = 4. A.3 B.5 1/3 C.18.75

OpenStudy (johnweldon1993):

Now this new question... again it varies directly so \[x = ky\] We know that x = 7.5 when y = 10 so \[7.5 = k(10)\] Divide both sides by 10 to solve for 'k' \[k = \frac{ 7.5 }{ 10 }\] Now that we know what 'k' equals...lets solve this problem... what is 'x' when y = 4? \[x = ky\] \[x = \frac{ 7.5 }{ 10 }(4)\] And we solve that to get...?

OpenStudy (anonymous):

C?

OpenStudy (johnweldon1993):

7.5 times 4 = 30 30/10 = ...?

OpenStudy (anonymous):

3

OpenStudy (johnweldon1993):

There you go

OpenStudy (anonymous):

: If t varies as v, and t = 2 4/7 when v =13/14, find v when t = 2 1/4. A.2106/392 B.13/16 C.324/52

OpenStudy (anonymous):

problem with fractions i kind stuck here

OpenStudy (johnweldon1993):

varies directly? or inversely?

OpenStudy (anonymous):

I'm not sure

OpenStudy (johnweldon1993):

I'm going to assume directly \[t = kv\] t = 2 4/7 when v =13/14 so *is that 24/7 or 2 4/7? fraction or mixed number?

OpenStudy (anonymous):

\[2 \frac{ 4 }{ 7 }\]

OpenStudy (johnweldon1993):

ahh okay a mixed number....convert that to an improper fraction for me...

OpenStudy (johnweldon1993):

Did you know how to?

OpenStudy (johnweldon1993):

So you take the denominator (7) and multiply it by the whole number (2)....and add the numerator to that..(4) so 7 times 2 = 14....+ 4 = 18....and you put this over the original denominator 18/7 is our improper fraction... so we have \[t = kv\] t = 18/7 when v =13/14 so \[\frac{ 18 }{ 7 } = k(\frac{ 13 }{ 14 })\] Now multiply both sides of the equation by 14/13 to isolate 'k' \[k = \frac{ 18 }{ 7 } \times \frac{ 14 }{ 13 }\] \[k = \frac{ 252 }{ 91 }\] Now that we know what 'k' equals..we can solve this

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

find v when t = 2 1/4. Turning that into an improper fraction we have 9/4 so \[t = kv\] \[\frac{ 9 }{ 4 } = \frac{ 252 }{ 91} v\] multiply both sides by 9/252 to isolate v... \[v = \frac{ 9 }{ 4 } \times \frac{ 91 }{ 252 }\] \[v = \frac{ 13 }{ 16 }\]

OpenStudy (anonymous):

:) thnx i will safe this work

OpenStudy (anonymous):

one sec i show one screen i have no clue how to solve this

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

you know how to do this?

OpenStudy (anonymous):

p ~ 1/v my answer but it doesn't match may be im wrong

OpenStudy (johnweldon1993):

that's weird....p = k/v is a proportion....it is inversely proportional...hmm I have no idea? Maybe try a direct proportion? pv = k? idk no idea on this one

OpenStudy (anonymous):

@johnweldon1993 one thing on screen they put that i just need to continue

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

so they put the K_1 = thing? try putting "pv"

OpenStudy (anonymous):

i can't remove

OpenStudy (johnweldon1993):

Yeah see I have no idea...

OpenStudy (anonymous):

i have to put whats on grid

OpenStudy (anonymous):

The volume (v) of a sphere varies directly as the cube of its diameter (d). Write this statement in algebraic language, using an equation with the variables c, v, and d.

OpenStudy (anonymous):

\[C=\frac{ ? }{ ? }\]

OpenStudy (anonymous):

same here they started C= ?/?

OpenStudy (anonymous):

@johnweldon1993 ?